OpenStudy (anonymous):

Which ratio is equal to 12:54 A. 3:13 B. 6:24 C. 24:104 D. 36:162

OpenStudy (tkhunny):

Hint: \(12:54 \implies \dfrac{12}{54} = \dfrac{6}{27} = \dfrac{2}{9}\)

OpenStudy (tkhunny):

A: 3:13 is already in it's lowest terms. So no good, there. B: \(6:24 \implies \dfrac{6}{24} = \dfrac{3}{12} = \dfrac{1}{4} \ne \dfrac{2}{9}\) Why did you think it might be 6:24? Seems a little like you were just guessing. Don't do that. Prove it!!

OpenStudy (anonymous):

i never understood ratios :(

OpenStudy (tkhunny):

Now is the time.

OpenStudy (anonymous):

well proportions ratios...

OpenStudy (tkhunny):

We're just hunting for common factors. \(6:24 \implies \dfrac{6}{24}\) -- You just have to know that. Both numerator and denominator are even. There must be a common factor of 2. \(\dfrac{6}{24} = \dfrac{2\cdot 3}{2\cdot 12} = \dfrac{2}{2}\cdot\dfrac{3}{12} = 1\cdot \dfrac{3}{12} = \dfrac{3}{12}\) Both numerator and denominator are divisible by 3. You show me how that works.

OpenStudy (anonymous):

dividing by 3s?

OpenStudy (tkhunny):

Perfect. That gives 1/4 as I demonstrated above. Okay, let's move on to C: \(24:104\implies \dfrac{24}{104}\) -- What to do first?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

by 24?

OpenStudy (tkhunny):

I don't think so. 104/24 = 4.333... Think smaller. Generally shoot for prime numbers. The easiest is to notice that they are both even. There's a 2 in there.

OpenStudy (anonymous):

so by 4rs?

OpenStudy (tkhunny):

4 is not a prime number, but after doing 2, we would have discovered that there is another 2. Thus, 4 is okay on this one. Go ahead and do it. Let's see what you get.

OpenStudy (anonymous):

thanks but i have to move on to my next question.

OpenStudy (tkhunny):

Ah,well... Come back when you're ready.

OpenStudy (anonymous):

its D

OpenStudy (tkhunny):

@mariruiz Please read the Code of Conduct and heed its wisdom.

OpenStudy (anonymous):

i was about to explain why

OpenStudy (tkhunny):

After the OP already signed off? A little odd, but okay.

OpenStudy (anonymous):

@Happie cross multiply (12*162) which gets you 1944. then multiply (36*54) which gets you 1944.

OpenStudy (anonymous):

@tkhunny i hadn't realized they signed off

OpenStudy (anonymous):

what? o.o

OpenStudy (anonymous):

its a shortcut to finding equivalent ratios @Happie

OpenStudy (anonymous):

the correct awnser was D. tkhunny u were way off.

OpenStudy (tkhunny):

?? That makes no sense. My purpose was to teach you how to find the answer on your own. If I was "way off", this can mean only that you only wanted the answer and didn't care to learn. I'm guessing that's not it, so I couldn't have been off at all. I ruled out A We ruled out B, together. We were working on C, and this was not yet done when you signed off. We hadn't yet gotten to D. I never use the profanity "cross-multiply". Do you want to learn it, or not?

OpenStudy (tkhunny):

(sigh)